# Infinitely many pairs of non-isomorphic elliptic curves sharing the same BSD invariants

### v1

@@ Abstract.
We prove that there exist infinitely many pairs of non-isomorphic elliptic curves over $\mathbb{Q}$ sharing the same BSD invariants—including their $L$-functions, Mordell–Weil groups, Tate–Shafarevich groups, Tamagawa numbers, regulators, and real periods—and Kodaira symbols and minimal discriminants, while having distinct $j$-invariants.
@@ 1. Introduction
Let $E/\mathbb{Q}$ be an elliptic curve defined over $\mathbb{Q}$. We define the following 6-tuple as Birch and Swinnerton-Dyer (BSD) data following $L$-functions and Modular Forms Database (LMFDB) [7]:
$\mathrm{BSD}(E/\mathbb{Q})\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\bigl{(}L(E,s),\,E(\mathbb{Q}),\,\mathrm{Reg}(E/\mathbb{Q}),\,\Omega_{E},\,c_{E/\mathbb{Q}_{p}},\,\Sha(E/\mathbb{Q})\bigr{)}.$
@@
When $E_{1}\cong E_{2}$ over $\mathbb{Q}$, we have $\text{BSD}(E_{1}/\mathbb{Q})=\text{BSD}(E_{2}/\mathbb{Q})$.
@@
In the high-dimensional case, Jamie Bell proved the existence of pairs of abelian varieties of dimension $p-1$ (where the class number of $\mathbb{Q}(\zeta_{p})$ is not $1$, so $p\geq 23$), defined over $\mathbb{Q}$, such that over every number field they share the same BSD invariants (except for the real period), the same $n$-Selmer groups for all $n$, and the same Tate modules [2]. The pair becomes isomorphic over $\overline{\mathbb{Q}}$. There are many coprime degree isogenies between this pair of abelian varieties. However, for two non-isomorphic elliptic curves over $\mathbb{Q}$, there cannot exist two or more isogenies of coprime degrees. Therefore, we need to examine the $q$-primary component of the invariants in the case of elliptic curves connected by prime degree $q$ isogenies.
@@
As noted in Remark 5.8, it has been proved that—up to twists—there is exactly one pair of 2‑isogenous elliptic curves that share the same BSD invariants, Kodaira symbols, and minimal discriminants. In summary, up to quadratic twist, there exists exactly one 2‑isogenous pair of elliptic curves over $\mathbf{Q}$ that cannot be distinguished by their BSD invariants, Kodaira symbols, minimal discriminants. And such indistinguishable pairs occur with positive density among quadratic twists of elliptic curves satisfying the sufficient condition of Proposition 5.4.
@@ 2.1. Notation
    $A[p]$ :=$:=\{\,a\in A\mid pa=0\}$ ($p$‑torsion subgroup)
    $A[p^{\infty}]$ :=$:=\{\,a\in A\mid\exists\,n\geq 1,\;p^{n}a=0\}$ ($p$‑primary component)
@@
  • For an isogeny $\phi$ between abelian varieties, $\hat{\phi}$ denotes the dual isogeny of $\phi$.
  • For an elliptic curve $E/\mathbb{Q}$, another elliptic curve $E^{\prime}/\mathbb{Q}$ and a nonzero isogeny $\phi:E\to E^{\prime}$, $\mathrm{Sel}^{\phi}(E/\mathbb{Q})$ is the $\phi$-Selmer group. When $E=E^{\prime}$ and $\phi=[n]$ (multiplication-by-$n\geq 1$ map), we denote its Selmer group by $\mathrm{Sel}^{n}(E/\mathbb{Q})$.
@@ 2.2. Outline of the proof
The LMFDB already contains several pairs satisfying this condition, and one can easily confirm that these pairs indeed have the same BSD invariants and Kodaira symbols by Proposition 5.4. Applying quadratic twists then yield infinitely many further pairs; the existence of suitable twists follows in general from the Smith density theorem or classical two descent (Proposition 6.1) in the special case that we have sought.
@@
where $\alpha_{r,u}$ denotes the limit, as $n\to\infty$, of the probability that a uniformly random $(n-u)\times n$ matrix over $\mathbb{F}_{2}$ has rank exactly $n-r$.
@@
where $W_{v}(\phi,D)\;:=\;\ker\mskip-3.0mu\Bigl{(}H^{1}(\mathbb{Q}_{v},E_{1}^{D}[\phi])\longrightarrow H^{1}(\mathbb{Q}_{v},E_{1}^{D})\Bigr{)}$  (see Theorem 3.5 of [6]).
When $E^{D}/\mathbb{Q}$ has good reduction at an odd prime $v$, $W_{v}(\phi,D)\cong H^{1}_{\mathrm{nr}}(\mathbb{Q}_{v},E^{D}[\phi])\cong H^{1}(\widehat{\mathbb{Z}},\mathbb{Z}/2\mathbb{Z})\cong\mathbb{Z}/2\mathbb{Z}$ where $H^{1}_{\mathrm{nr}}(\mathbb{Q}_{v},E^{D}[\phi])$ is the unramified cohomology. When $E/\mathbb{Q}$ has good reduction at an odd prime $v$ and $v(D)=1$, $W_{v}(\phi,D)\cong\mathbb{Z}/2\mathbb{Z}$ (see section 3.1 of [11]). The remaining cases are where $E/\mathbb{Q}$ has bad reduction at $v$ or $v=2$ or $v=\infty$. Since $DD_{0}\in(\mathbb{Q}_{v}^{\times})^{2}$ for every $v\in S$, $E_{1}^{D}[\phi]\cong E_{1}^{D_{0}}[\phi]$ and $W_{v}(\phi,D)\cong W_{v}(\phi,D_{0})$. Summarizing the above, $\tau^{D}=\tau^{D_{0}}$. We use this remark to simultaneously control $\Sha(E_{1}^{D}/\mathbb{Q})$ and $\Sha(E_{2}^{D}/\mathbb{Q})$ in Section 6.
@@
By Theorem 3.2, there exist infinitely many square-free integers $D$ such that $(D,2\Delta_{E_{1}}\Delta_{E_{2}})=1$ and $e_{D}=f_{D}=1$ by taking $D_{0}=1$ and $r=0$. Indeed, with $D_{0}=1$, we have $u=e_{D}-f_{D}=\mathrm{dim}_{\mathbb{F}_{2}}\mathrm{Sel}^{\phi}(E_{1}^{D}/\mathbb{Q})-\mathrm{dim}_{\mathbb{F}_{2}}\mathrm{Sel}^{\hat{\phi}}(E_{2}^{D}/\mathbb{Q})=0$. Also, $D\in(\mathbb{Q}_{v}^{\times})^{2}$ for every $p\mid 2\Delta_{E_{1}}\Delta_{E_{2}}$ implies $\text{gcd}(D,2\Delta_{E_{1}}\Delta_{E_{2}})=1$. Hence, it remains only to show that $\alpha_{0,0}>0$. In fact,
@@
(3.5) #$\#\Sha(E^{D}_{i}/\mathbb{Q})[2]=\dim_{\mathbb{F}_{2}}\mathrm{Sel}^{2}(E^{D}_{i}/\mathbb{Q})-\dim_{\mathbb{F}_{2}}E^{D}_{i}(\mathbb{Q})/2E^{D}_{i}(\mathbb{Q})\leq 1-1=0$
for $i\in\{1,2\}$. Additionally, since $\dfrac{E^{D}_{i}(\mathbb{Q})}{2E^{D}_{i}(\mathbb{Q})}\leq\mathrm{Sel}^{2}(E^{D}_{i}/\mathbb{Q})\leq 2$ and #$\#\dfrac{E^{D}_{i}(\mathbb{Q})}{2E^{D}_{i}(\mathbb{Q})}=2^{\mathrm{rank}(E^{D}_{i}/\mathbb{Q})+1}$, we have $\mathrm{rank}(E^{D}_{i}/\mathbb{Q})=0$ for $i\in\{1,2\}$.
@@ 4. Kodaira symbol, Tamagawa number and real period
Let $E_{1}$ and $E_{2}$ be isogenous elliptic curves over $\mathbb{Q}$, and assume $E_{1}$ has good reduction at $2$. The following are equivalent:
@@
($(1)\Rightarrow(2)$) Suppose all Kodaira symbols are the same. Ogg’s formula states that
$v_{l}(\Delta_{E})=f_{l}+m_{l}-1,$
where $f_{l}$ is the conductor exponent and $m_{l}=\#\tilde{\epsilon}(\overline{\mathbb{F}_{l}})/\#\tilde{\epsilon}_{0}(\overline{\mathbb{F}_{l}})$. Here, $\tilde{\epsilon}$ is the special fiber of the Néron minimal model of $E/\mathbb{Q}$, $\tilde{\epsilon}_{0}$ is its connected component of the identity, and $\Delta_{E}$ is the minimal discriminant of $E/\mathbb{Q}$. Since $E_{1}$ and $E_{2}$ are isogenous, the conductor exponent $f_{l}$ is an isogeny invariant, so $f_{l,E_{1}}=f_{l,E_{2}}$ for all primes $l$. Since $E_{1}$ and $E_{2}$ have the same Kodaira symbols at all primes $l$, we have $m_{l,E_{1}}=m_{l,E_{2}}$ for all primes $l$. Therefore, $v_{l}(\Delta_{E_{1}})=v_{l}(\Delta_{E_{2}})$ for all primes $l$. This implies $\Delta_{E_{1}}=\pm\Delta_{E_{2}}$.
($(1)\Leftarrow(2)$) Suppose $\Delta_{E_{1}}=\pm\Delta_{E_{2}}$. Then for all primes $\ell$, we have $v_{\ell}(\Delta_{E_{1}})=v_{\ell}(\Delta_{E_{2}})$. By Ogg’s formula, this means $f_{\ell,E_{1}}+m_{\ell,E_{1}}-1=f_{\ell,E_{2}}+m_{\ell,E_{2}}-1$. Since $E_{1}$ and $E_{2}$ are isogenous, $f_{\ell,E_{1}}=f_{\ell,E_{2}}$, which implies $m_{\ell,E_{1}}=m_{\ell,E_{2}}$. By lemma 2.16 of [1], $E_{1}$ has potentially good reduction at $l>2$. In the additive potentially good case, the Kodaira symbol is uniquely determined by $m_{\ell}$ (See [10], IV.9, Table 4.1). Since both invariants are equal, the Kodaira symbols are the same. ∎
@@
Suppose $E_{1}/\mathbb{Q}$ and $E_{2}/\mathbb{Q}$ are elliptic curves such that there exists a degree $2$ isogeny between them. Let $l\geq 3$ be a prime and $F$ be a quadratic extension of $\mathbb{Q}_{l}$ over which $E_{1}$ (or $E_{2}$) has good reduction. If $\frac{\Delta_{E_{1}}}{\Delta_{E_{2}}}\in N_{F/\mathbb{Q}_{l}}$, then $c_{E_{1}/\mathbb{Q}_{l}}=c_{E_{2}/\mathbb{Q}_{l}}$.
@@
See [3], Table 1 and Theorem 6.1. ∎
Proposition 4.3.
Let $p$ be a prime number. Let $E_{1}/\mathbb{Q}$ and $E_{2}/\mathbb{Q}$ be elliptic curves that are isogenous by a degree $p$ isogeny over $\mathbb{Q}$. Assume that $E_{1}$ and $E_{2}$ have the same Tamagawa numbers at all primes and $\mathrm{rank}_{an}(E_{1}/\mathbb{Q})\equiv 0\bmod 2$. Then, $\Omega_{E_{1}}=\Omega_{E_{2}}$.
In particular, if $\text{BSD}(E_{1}/\mathbb{Q})=\text{BSD}(E_{2}/\mathbb{Q})$ and $E_{1}$ and $E_{2}$ are isogenous by degree $p$ isogeny, then $E_{1}/\mathbb{Q}$ and $E_{2}/\mathbb{Q}$ have even analytic rank.
@@
$\Omega_{E_{1}}=\Omega_{E_{2}}\Leftrightarrow\sum_{l}\text{ord}_{p}(\frac{c_{E_{1}/\mathbb{Q}_{l}}}{c_{E_{2}/\mathbb{Q}_{l}}})\equiv\text{ord}_{s=1}L(E_{1},s)\bmod 2.$
Since Tamagawa numbers at all primes are the same, $\sum_{l}\text{ord}_{p}(\frac{c_{E_{1}/\mathbb{Q}_{l}}}{c_{E_{2}/\mathbb{Q}_{l}}})=0$ holds. Thus, the equality $\Omega_{E_{1}}=\Omega_{E_{2}}$ is equivalent to the congruence $\text{ord}_{s=1}L(E_{1},s)\equiv 0\bmod 2$, which is true.

### v2

@@ Abstract.
We prove that there exist infinitely many pairs of non-isomorphic elliptic curves over $\mathbb{Q}$ sharing the same BSD invariants—including their Mordell–Weil groups, Tate–Shafarevich groups, Tamagawa numbers, regulators, and real periods—and Kodaira symbols and minimal discriminants, while having distinct $j$-invariants.
@@ 1. Introduction
Let $E/\mathbb{Q}$ be an elliptic curve defined over $\mathbb{Q}$. We define the following 6-tuple as Birch and Swinnerton-Dyer (BSD) data:
$\mathrm{BSD}(E/\mathbb{Q})\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\bigl{(}L(E,s),\,E(\mathbb{Q}),\,\mathrm{Reg}(E/\mathbb{Q}),\,\Omega_{E},\,(c_{E/\mathbb{Q}_{p}})_{p},\,\Sha(E/\mathbb{Q})\bigr{)}.$
@@
When $E_{1}\cong E_{2}$ over $\mathbb{Q}$, we have $\text{BSD}(E_{1}/\mathbb{Q})=\text{BSD}(E_{2}/\mathbb{Q})$. Here, “=” between groups means that the groups are isomorphic as groups.
@@
In the high-dimensional case, Jamie Bell proved the existence of pairs of abelian varieties of dimension p1p-1 (where the class number of $\mathbb{Q}(\zeta_{p})$ is not $1$, so $p\geq 23$), defined over $\mathbb{Q}$, such that over every number field they share the same BSD invariants (except for the real period), the same $n$-Selmer groups for all $n$, and the same Tate modules [2]. The pair becomes isomorphic over $\overline{\mathbb{Q}}$. There are many coprime degree isogenies between this pair of abelian varieties. However, for two non-isomorphic elliptic curves over $\mathbb{Q}$, there cannot exist two or more isogenies of coprime degrees. Therefore, we need to examine the $q$-primary component of the invariants in the case of elliptic curves connected by prime degree $q$ isogenies.
@@
As noted in Remark 5.8, it has been proved that—up to twists—there is a unique pair of 2‑isogenous elliptic curves that share the same BSD invariants, Kodaira symbols, and minimal discriminants. In summary, up to quadratic twist, there exists exactly one 2‑isogenous pair of elliptic curves over $\mathbb{Q}$ that cannot be distinguished by their BSD invariants, Kodaira symbols, minimal discriminants. And such indistinguishable pairs occur with positive density among quadratic twists of elliptic curves satisfying the sufficient condition of Proposition 5.4.
@@ 2.1. Notation
    $A[p]$ $=\{\,a\in A\mid pa=0\}$ (pp‑torsion subgroup)
    $A[p^{\infty}]$ $=\{\,a\in A\mid\exists\,n\geq 1,\;p^{n}a=0\}$ (pp‑primary component)
@@
  • For an isogeny $\phi:E_{1}\to E_{2}$ between elliptic curves, $\hat{\phi}:E_{2}\to E_{1}$ denotes the dual isogeny of $\phi$.
  • For a quadratic twist $E^{D}$ of an elliptic curve $E$ and a degree $2$ isogeny $\phi:E_{1}\to E_{2}$ defined over $\mathbb{Q}$, we also denote by $\phi$ the induced degree $2$ isogeny $\phi_{D}:E_{1}^{D}\to E_{2}^{D}$ defined over $\mathbb{Q}$. Accordingly, we abbreviate the $\phi_{D}$-Selmer group $\operatorname{Sel}^{\phi_{D}}(E_{1}^{D}/\mathbb{Q})$ to $\operatorname{Sel}^{\phi}(E_{1}^{D}/\mathbb{Q})$.
@@ 2.2. Outline of the proof
The LMFDB already contains several pairs satisfying this condition, and one can easily confirm that these pairs indeed have the same BSD invariants and Kodaira symbols by Proposition 5.4. Applying quadratic twists then yields infinitely many further pairs; the existence of suitable twists follows in general from the Smith density theorem or classical two descent (Proposition 6.1) in the special case that we have sought.
@@
where $\alpha_{r,u}$ denotes the limit, as $n\to\infty$, of the probability that a uniformly random $(n-u)\times n$ matrix over $\mathbb{F}_{2}$ has rank exactly nrn-r.
@@
where $W_{v}(\phi,D)\;:=\;\text{Ker}\ \mskip-3.0mu\Bigl{(}H^{1}(\mathbb{Q}_{v},E_{1}^{D}[\phi])\longrightarrow H^{1}(\mathbb{Q}_{v},E_{1}^{D})\Bigr{)}$  (see Theorem 3.5 of [6]).
When $E^{D}/\mathbb{Q}$ has good reduction at an odd prime $v$, $W_{v}(\phi,D)\cong H^{1}_{\mathrm{nr}}(\mathbb{Q}_{v},E^{D}[\phi])\cong H^{1}(\widehat{\mathbb{Z}},\mathbb{Z}/2\mathbb{Z})\cong\mathbb{Z}/2\mathbb{Z}$ where $H^{1}_{\mathrm{nr}}(\mathbb{Q}_{v},E^{D}[\phi])$ is the unramified cohomology. When $E/\mathbb{Q}$ has good reduction at an odd prime $v$ and $v(D)=1$, $W_{v}(\phi,D)\cong\mathbb{Z}/2\mathbb{Z}$ (see section 3.1 of [12]). The remaining cases are where $E/\mathbb{Q}$ has bad reduction at $v$ or $v=2$ or $v=\infty$. Since $DD_{0}\in(\mathbb{Q}_{v}^{\times})^{2}$ for every $v\in S$, $E_{1}^{D}[\phi]\cong E_{1}^{D_{0}}[\phi]$ and $W_{v}(\phi,D)\cong W_{v}(\phi,D_{0})$. Summarizing the above, $\tau^{D}=\tau^{D_{0}}$. We use this remark to simultaneously control $\Sha(E_{1}^{D}/\mathbb{Q})$ and $\Sha(E_{2}^{D}/\mathbb{Q})$ in the following Proposition.
@@
By Theorem 3.2, there exist infinitely many square-free integers $D$ such that $(D,2\Delta_{E_{1}}\Delta_{E_{2}})=1$ and $e_{D}=f_{D}=1$ by taking $D_{0}=1$ and $r=0$. Indeed, with $D_{0}=1$, we have $u=e_{D}-f_{D}=\mathrm{dim}_{\mathbb{F}_{2}}\mathrm{Sel}^{\phi}(E_{1}^{D}/\mathbb{Q})-\mathrm{dim}_{\mathbb{F}_{2}}\mathrm{Sel}^{\hat{\phi}}(E_{2}^{D}/\mathbb{Q})=\mathrm{dim}_{\mathbb{F}_{2}}\mathrm{Sel}^{\phi}(E_{1}/\mathbb{Q})-\mathrm{dim}_{\mathbb{F}_{2}}\mathrm{Sel}^{\hat{\phi}}(E_{2}/\mathbb{Q})=0$ by Remark 3.3. Also, $D\in(\mathbb{Q}_{v}^{\times})^{2}$ for every $p\mid 2\Delta_{E_{1}}\Delta_{E_{2}}$ implies $\text{gcd}(D,2\Delta_{E_{1}}\Delta_{E_{2}})=1$. Hence, it remains only to show that $\alpha_{0,0}>0$. In fact,
@@
(3.5) $\dim_{\mathbb{F}_{2}}\Sha(E^{D}_{i}/\mathbb{Q})[2]=\dim_{\mathbb{F}_{2}}\mathrm{Sel}^{2}(E^{D}_{i}/\mathbb{Q})-\dim_{\mathbb{F}_{2}}E^{D}_{i}(\mathbb{Q})/2E^{D}_{i}(\mathbb{Q})\leq 1-1=0$
for $i\in\{1,2\}$. Additionally, since $\dfrac{E^{D}_{i}(\mathbb{Q})}{2E^{D}_{i}(\mathbb{Q})}\leq\#\mathrm{Sel}^{2}(E^{D}_{i}/\mathbb{Q})\leq 2$ and $\dfrac{E^{D}_{i}(\mathbb{Q})}{2E^{D}_{i}(\mathbb{Q})}=2^{\mathrm{rank}(E^{D}_{i}/\mathbb{Q})+1}$, we have $\mathrm{rank}(E^{D}_{i}/\mathbb{Q})=0$ for $i\in\{1,2\}$.
@@ 4. Kodaira symbol, Tamagawa number and real period
Let $p$ be a prime number. Let $E_{1}/\mathbb{Q}$ and $E_{2}/\mathbb{Q}$ be elliptic curves that are isogenous by a degree $p$ isogeny over $\mathbb{Q}$. Assume $E_{1}$ has good reduction at $2$. The following are equivalent:
@@
($(1)\Rightarrow(2)$) Ogg’s formula states that
$v_{\ell}(\Delta_{E})=f_{\ell,E}+m_{\ell,E}-1,$
where $f_{\ell}$ is the conductor exponent and $m_{\ell}=\#(\tilde{\epsilon}/\tilde{\epsilon}_{0})(\overline{\mathbb{F}_{\ell}})$(see [10], §11). Here, $\tilde{\epsilon}$ is the special fiber of the Néron minimal model of $E/\mathbb{Q}$, $\tilde{\epsilon}_{0}$ is its connected component of the identity, and $\Delta_{E}$ is the minimal discriminant of $E/\mathbb{Q}$. Since $E_{1}$ and $E_{2}$ are isogenous, the conductor exponent $f_{\ell}$ is an isogeny invariant, so $f_{\ell,E_{1}}=f_{\ell,E_{2}}$ for all primes $\ell$. Since $E_{1}$ and $E_{2}$ have the same Kodaira symbols at all primes $\ell$, we have $m_{\ell,E_{1}}=m_{\ell,E_{2}}$ for all primes $\ell$. Therefore, $v_{\ell}(\Delta_{E_{1}})=v_{\ell}(\Delta_{E_{2}})$ for all primes $\ell$. This implies $\Delta_{E_{1}}=\pm\Delta_{E_{2}}$.
($(1)\Leftarrow(2)$) Suppose $\Delta_{E_{1}}=\pm\Delta_{E_{2}}$. Then for all primes $\ell$, we have $v_{\ell}(\Delta_{E_{1}})=v_{\ell}(\Delta_{E_{2}})$. By Ogg’s formula, this means $f_{\ell,E_{1}}+m_{\ell,E_{1}}-1=f_{\ell,E_{2}}+m_{\ell,E_{2}}-1$. Since $E_{1}$ and $E_{2}$ are isogenous, $f_{\ell,E_{1}}=f_{\ell,E_{2}}$, which implies $m_{\ell,E_{1}}=m_{\ell,E_{2}}$. By Lemma 2.16 of [1], $E_{1}$ has potentially good reduction at $\ell$. In the additive potentially good case, the Kodaira symbol at $\ell>2$ is uniquely determined by $m_{\ell}$ (See [11], IV.9, Table 4.1). Since $m_{\ell,E_{1}}=m_{\ell,E_{2}}$, the Kodaira symbols are the same. ∎
@@
Suppose $E_{1}/\mathbb{Q}$ and $E_{2}/\mathbb{Q}$ are elliptic curves such that there exists a degree $2$ isogeny defined over $\mathbb{Q}$ between them. Let $\ell\geq 3$ be a prime. If $\Delta_{E_{1}}=\Delta_{E_{2}}$, then $c_{E_{1}/\mathbb{Q}_{\ell}}=c_{E_{2}/\mathbb{Q}_{\ell}}$.
@@
Since $\Delta_{E_{1}}=\Delta_{E_{2}}$, $E_{1}$ and $E_{2}$ have potentially good reduction at $\ell$ by Lemma 2.16 of [1]. When the Kodaira symbol at $\ell$ is not $I^{*}_{0}$, then $c_{E_{1}/\mathbb{Q}_{\ell}}=c_{E_{2}/\mathbb{Q}_{\ell}}$ (see the proof of Theorem 6.1, [3]). When the Kodaira symbol at $\ell$ is $I^{*}_{0}$, the result follows since $1=\frac{\Delta_{E_{1}}}{\Delta_{E_{2}}}\in N_{F/\mathbb{Q}_{\ell}}$ where $F$ is a quadratic extension of $\mathbb{Q}_{\ell}$ over which $E_{1}$ has good reduction (see [3], Table 1 and Theorem 6.1).
Theorem 4.3.
Let $p$ be a prime number. Let $E_{1}/\mathbb{Q}$ and $E_{2}/\mathbb{Q}$ be elliptic curves that are isogenous by a degree $p$ isogeny defined over $\mathbb{Q}$. Assume that $E_{1}$ and $E_{2}$ have the same Tamagawa numbers at all primes and $\mathrm{rank}_{an}(E_{1}/\mathbb{Q})\equiv 0\bmod 2$. Then, $\Omega_{E_{1}}=\Omega_{E_{2}}$.
In particular, if $\text{BSD}(E_{1}/\mathbb{Q})=\text{BSD}(E_{2}/\mathbb{Q})$ and $E_{1}$ and $E_{2}$ are isogenous by degree $p$ isogeny defined over $\mathbb{Q}$, then $E_{1}/\mathbb{Q}$ and $E_{2}/\mathbb{Q}$ have even analytic rank.
@@
$\Omega_{E_{1}}=\Omega_{E_{2}}\Leftrightarrow\sum_{\ell}\text{ord}_{p}(\frac{c_{E_{1}/\mathbb{Q}_{\ell}}}{c_{E_{2}/\mathbb{Q}_{\ell}}})\equiv\text{ord}_{s=1}L(E_{1},s)\bmod 2.$
Since Tamagawa numbers at all primes are the same, $\sum_{l}\text{ord}_{p}(\frac{c_{E_{1}/\mathbb{Q}_{\ell}}}{c_{E_{2}/\mathbb{Q}_{\ell}}})=0$ holds. Thus, the equality $\Omega_{E_{1}}=\Omega_{E_{2}}$ is equivalent to the congruence $\text{ord}_{s=1}L(E_{1},s)\equiv 0\bmod 2$, which is true.